The accounting records of Wall's China Shop reflected the following balances as of January 1, Year 2: Cash $ 18,900 Beginning inventory: 14,105 (155 units @ $91) Common stock: 16,000 Retained earnings: 17,005 The following five transactions occurred in Year 2: First purchase (cash) 120 units @ $93 Second purchase (cash) 200 units @ $101 Sales (all cash) 420 units @ $191 Paid $15,100 cash for salaries expenses. Paid cash for income tax at the rate of 40 percent of income before taxes. Required: a. Compute the cost of goods sold and ending inventory, assuming (1) FIFO cost flow, (2) LIFO cost flow, and (3) weighted-average cost flow. Compute the income tax expense for each method. b. (this part has been omitted) c. Use a vertical model to show the Year 2 income statement, balance sheet, and statement of cash flows under FIFO, LIFO, and weighted average.
Answer #1
Opening Trial balance:
Debit Credit
Cash 18900
Inventory (155@91) 14105
Common Stock 16000
Retained earnings                    17005
Total 33005 33005
Inventory account:
Opening balance 155 @ 91 14105
Purchases:
Cash 120 @ 93 11160
200 @ 101 20200
Total purchases 31360
Total inventory 475 45465
Average cost 95.71579
Sales 420
Closing Inventory 55 units
Expenses (salary) 15100
a) COGS:
FIFO = 14105+11160+(145*101)=39910
LIFO= 20200+11160+(100*91)=40460
WA=420*95.71579 = 40201
Ending inventory:
FIFO=55*101=5555
LIFO=55*91=5005
WA=55*95.71579=5264
c)
Income Statement:
FIFO LIFO WA
Sales 80220 80220 80220 420*191
Less:COGS 39910 40460 40201
Gross Profit 40310 39760 40019
Less:Expenses, salary 15100 15100 15100
IBT 25210 24660 24919
Less : Income Tax@40% 10084 9864 9968
Net Income 15126 14796 14951
Balance Sheet:
Assets
Cash 42576 42796 42692 18900+80220-31360-15100-10084
Inventory 5555 5005 5264
Total Assets 48131 47801 47956
Liabilities
Common Stock 16000 16000 16000
Retained Earnings 32131 31801 31956 17005+15126
Total Liabilities 48131 47801 47956
Cash Flow :
Net Income 15126 14796 14951
add:decrease in Inventory 8550 9100 8841 14105-5555
Net Cash inflow 23676 23896 23792
add: Opening Cash 18900 18900 18900
Closing Cash 42576 42796 42692